A get 20% more money than B.how much% money B gets less than A???? sir plllzzz help to solve this type of questions plzzz sir as soon as posibl
Answer by jim_thompson5910(35256)

Let's say B earns $100

A would earn $120 since 20% of 100 is 20. Then you add that onto 100

-------------------------------------------------------

120 - 100 = 20

Divide 20 by 120 to get 20/120 = 2/12 = 1/6 = 0.1667 (roughly)

So B learns roughly 16.67% less than A